# Kino-PAX: Highly Parallel Kinodynamic Sampling-based Planner

### v1

@@ IV-B Efficient Application to Parallel Devices
Furthermore, the organization of samples into three disjoint sets, $V_{U},V_{O},$ and $V_{E}$, enables a straightforward and memory-efficient representation. In Kino-PAX, we do this via a boolean-mask representation that is thoroughly discussed in [29]. For completeness, we also briefly describe the implementation below.
Kino-PAX’s three sets partition all samples through boolean flags, where a sample (i.e., an $n$-dimensional floating point representation) is stored in a one dimensional $t_{e}$-sized array, and its corresponding index in the large array is set to $true$ in one of the sets. We then efficiently access the correct sample information via unique thread identifiers and a prefix-sum operation over the boolean set which essentially equates to a parallelized graph search.
Finally, we pre-allocate a large memory chunk on the GPU that can accommodate $t_{e}$ nodes. This memory acts as an empty tree structure that is populated as planning progresses, eliminating the need for expensive vector resizing operations and allowing Kino-PAX to quickly iterate through the planning process. Additionally, this implementation keeps all large data structures entirely on the GPU. During planning, only small pieces of data are transferred to the CPU, reducing idle time between subroutine calls.
@@ V-A Probabilistic Completeness
Case 2: a region $\mathcal{R}_{i}$ is entirely within free space, i.e., $\mathcal{R}_{i}\subseteq X_{\text{free}}$. If $\mathcal{T}$ does not contain a node in $\mathcal{R}_{i}$, then $\mathcal{R}_{i}\notin\mathcal{R}_{avail}$, and (5) is set to $1>\epsilon$, as by Line 3 of Alg. 1. However, if $\mathcal{T}$ does contain a node in $\mathcal{R}_{i}$, we examine the worst-case scenario that produces the lowest probability from (5). In this scenario, the number of nodes in $\mathcal{R}_{i}$ approaches infinity, driving the score of $\mathcal{R}_{i}$ to 00 by (4). Additionally, in a worst-case scenario, the score of all other available regions approaches infinity. In this case, (5) trivially approaches $\epsilon$.
Case 3: $\mathcal{R}_{i}$ is partially in $X_{\text{free}}$, i.e., $\mathcal{R}_{i}\cap X_{\text{free}}\neq\emptyset$ and $\mathcal{R}_{i}\cap X_{\text{free}}\neq\mathcal{R}_{i}$. In this case, similar to above, if $\mathcal{R}_{i}\notin\mathcal{R}_{avail}$, (5) is set to $1$ by Line 3 of Alg. 1. If $\mathcal{R}_{i}\in\mathcal{R}_{avail}$, as in Case 2, (5) approaches $\epsilon$ in the worst-case, when the number of nodes in $\mathcal{R}_{i}$ approaches infinity. ∎
@@
Our proof is an adaptation of the result from [30], originally established for kinodynamic RRT, for Kino-PAX. Specifically, the conditions laid out in [30, Lemma 2–3] hold under the assumptions of Problem 1 in this paper. Furthermore, Lemma 1 provides a lower bound $\epsilon>0$ on the probability of extending from an arbitrary tree node. By combining [30, Lemma 2-3] with Lemma 1, we can follow the same logical structure as the proof of [30, Theorem 2]. Thus, as the number of samples approaches infinity, Kino-PAX asymptotically almost-surely finds a valid trajectory from $x_{\text{init}}$ to $X_{\text{goal}}$ that solves Problem 1. ∎
@@ VI Experiments
Table I: Benchmark results over 50 trials with maximum planning time of 60 seconds.
Algorithm Device            Environment 1            Environment 2            Environment 3
@@ VI-A Benchmark Results
For both 6D systems, Kino-PAX finds a solution trajectory in less than $8$ ms across all testing environments. For the 6D Double Integrator, Kino-PAX is on average 85×85$\times$85 ×, 287×287$\times$287 ×, 433×433$\times$433 × faster in Environments 1, 2, 3, respectively, compared to the baseline algorithms. For the Dubins Airplane system, the performance gap of Kino-PAX widens further. For instance, in Environment 3, Kino-PAX experiences a slowdown of less than 1.3×1.3$\times$1.3 × ($\sim$2 ms) compared to RRT’s $\times2$ ($\sim$20,000 ms), EST’s 1.8×1.8$\times$1.8 × ($\sim$4,000 ms), PDST’s 8.9×8.9$\times$8.9 × ($\sim$16,000 ms), and SyCLoP’s $\times2$ ($\sim$24,000 ms). Additionally, the embedded GPU implementation of Kino-PAX outperforms all serial baseline methods, finding valid trajectories for all 6D problems in under $115$ ms. When dealing with the more challenging 12D nonlinear quadcopter problem, Kino-PAX finds solutions in less than $25$ ms across all environments. On average, this marks an improvement of three orders of magnitude over all reference serial solutions. In the most challenging environment (Environment 3), the best-performing baseline algorithm (EST) is 1720×1720$\times$1720 × slower than the desktop implementation of Kino-PAX and $\times4$ slower than the embedded GPU implementation.
@@ VI-B Effects Of Tuning Parameter $t_{e}$
To support the point made in Sec. IV that Kino-PAX’s hyperparameter $t_{e}$ is easy to tune and that Kino-PAX remains efficient across a wide range of values, we present a numerical experiment showing that as Kino-PAX is provided with a sufficiently large $t_{e}$, its failure rate converges to zero. We also examine the impact on runtime as $t_{e}$ increases. Fig. 4 presents the results for planning with the 12D nonlinear quadcopter system in Environment 1.
@@
Figure 4: Results of varying the expected tree size $t_{e}$ for the 12D nonlinear quadcopter system in environment 3(a). (a) Number of Failures vs. $t_{e}$. (b) Mean runtime and variance of Kino-PAX vs. $t_{e}$.

### v2

@@ IV-C Efficient Application to Highly Parallel Devices
Furthermore, the organization of samples into three disjoint sets, $V_{U},V_{O},$ and $V_{E}$, enables a straightforward and memory-efficient representation. In Kino-PAX, we do this via a boolean-mask representation that is thoroughly discussed in [30].
Lastly, Kino-PAX reduces latency between its subroutines by pre-allocating a large memory chunk on the GPU to accommodate $t_{e}$ nodes. This allows Kino-PAX to construct its tree directly on the GPU, avoiding the transfer of large data structures between devices.
@@ V-A Probabilistic Completeness
Case 2: a region $\mathcal{R}_{i}$ is entirely within free space, i.e., $\mathcal{R}_{i}\subseteq X_{\text{valid}}$. If $\mathcal{T}$ does not contain a node in $\mathcal{R}_{i}$, then $\mathcal{R}_{i}\notin\mathcal{R}_{avail}$, and (5) is set to $1>\epsilon$, as by Line 3 of Alg. 1. However, if $\mathcal{T}$ does contain a node in $\mathcal{R}_{i}$, we examine the worst-case scenario that produces the lowest probability from (5). In this scenario, the number of nodes in $\mathcal{R}_{i}$ approaches infinity, driving the score of $\mathcal{R}_{i}$ to 00 by (4). Additionally, in a worst-case scenario, the score of all other available regions approaches infinity. In this case, (5) trivially approaches $\epsilon$.
Case 3: $\mathcal{R}_{i}$ is partially in $X_{\text{valid}}$, i.e., $\mathcal{R}_{i}\cap X_{\text{valid}}\neq\emptyset$ and $\mathcal{R}_{i}\cap X_{\text{valid}}\neq\mathcal{R}_{i}$. In this case, similar to above, if $\mathcal{R}_{i}\notin\mathcal{R}_{avail}$, (5) is set to $1$ by Line 3 of Alg. 1. If $\mathcal{R}_{i}\in\mathcal{R}_{avail}$, as in Case 2, (5) approaches $\epsilon$ in the worst-case, when the number of nodes in $\mathcal{R}_{i}$ approaches infinity. ∎
@@
Our proof is an adaptation of the result from [31], originally established for kinodynamic RRT, for Kino-PAX. Specifically, the conditions laid out in [31, Lemma 2–3] hold under the assumptions of Problem 1 in this paper. Furthermore, Lemma 1 provides a lower bound $\epsilon>0$ on the probability of extending from an arbitrary tree node. By combining [31, Lemma 2-3] with Lemma 1, we can follow the same logical structure as the proof of [31, Theorem 2]. Thus, as the number of samples approaches infinity, Kino-PAX asymptotically almost-surely finds a valid trajectory from $x_{\text{init}}$ to $X_{\text{goal}}$ that solves Problem 1. ∎
@@ VI Experiments
Table I: Benchmark results over 50 trials with a 60-second maximum planning time. CPU-based algorithms used coarse-grained parallelization, growing multiple trees in parallel and utilizing all available cores, denoted by “Par” before the algorithm name.
Algorithm Device            Environment LABEL:sub@fig:trees            Environment LABEL:sub@fig:narrowPassage            Environment LABEL:sub@fig:house
@@ VI-A Benchmark Results
For both 6D systems, Kino-PAX finds a solution trajectory in less than $8$ ms across all testing environments. For the 6D Double Integrator, Kino-PAX is on average 85×85$\times$85 ×, 287×287$\times$287 ×, 433×433$\times$433 × faster in Environments LABEL:sub@fig:trees, LABEL:sub@fig:narrowPassage, LABEL:sub@fig:house, respectively, compared to the baseline algorithms. For the Dubins Airplane system, the performance gap of Kino-PAX widens further. For instance, in Environment LABEL:sub@fig:house, Kino-PAX experiences a slowdown of less than 1.3×1.3$\times$1.3 × ($\sim$2 ms) compared to RRT’s $\times2$ ($\sim$20,000 ms), EST’s 1.8×1.8$\times$1.8 × ($\sim$4,000 ms), PDST’s 8.9×8.9$\times$8.9 × ($\sim$16,000 ms), and SyCLoP’s $\times2$ ($\sim$24,000 ms). Additionally, the embedded GPU implementation of Kino-PAX outperforms all serial baseline methods, finding valid trajectories for all 6D problems in under $115$ ms. When dealing with the more challenging 12D nonlinear quadcopter problem, Kino-PAX finds solutions in less than $25$ ms across all environments. On average, this marks an improvement of three orders of magnitude over all reference serial solutions. In the most challenging environment (Environment LABEL:sub@fig:house), the best-performing baseline algorithm (EST) is 1720×1720$\times$1720 × slower than the desktop implementation of Kino-PAX and $\times4$ slower than the embedded GPU implementation.
@@ VI-B Effects Of Tuning Parameter $t_{e}$
To support the point made in Sec. IV that Kino-PAX’s hyperparameter $t_{e}$ is easy to tune and that Kino-PAX remains efficient across a wide range of values, we present a numerical experiment showing that as Kino-PAX is provided with a sufficiently large $t_{e}$, its failure rate converges to zero. We also examine the impact on runtime as $t_{e}$ increases. Fig. 3 presents the results for planning with the 12D nonlinear quadcopter system in Environment LABEL:sub@fig:trees.
@@
Figure 3: Results of varying the expected tree size $t_{e}$ for the 12D nonlinear quadcopter system in environment 2(a). (a) Number of Failures vs. $t_{e}$. (b) Mean runtime and variance of Kino-PAX vs. $t_{e}$.